OpenStudy (anonymous):

(b+5)(b+10)=6 Help me solve please x= -11, x= -4 But how? Do you do it?

OpenStudy (anonymous):

i know it

OpenStudy (anonymous):

Yesss please do tell :)

OpenStudy (anonymous):

there is an identity that we can use to solve this .(x+a)(x+b)=x^2 +(a+b)x +ab.use this

OpenStudy (anonymous):

still don't get it :( how would i plug that in?

OpenStudy (anonymous):

x=b a =5 b=10

OpenStudy (anonymous):

(b+5)(b+10)=6^2 + (5+10)b+(5)(10) ?

OpenStudy (anonymous):

thank youuuu

myininaya (myininaya):

(b+5)(b+10)=6 b^2+10b+5b+50=6 b^2+15b+50=6 b^2+15b+50-6=0 b^2+15b+44=0 (b+11)(b+4)=0 b=-11 b=-4